\section*{Bibliography}
\begin{thebibliography}{00}

\bibitem{AlSo} V. Alexiades,  A.D. Solomon, Mathematical Modelling of Melting and Freezing Processes, Hemisphere-Taylor: Francis, Washington, 1993.

\bibitem{BeTaVi2000}	L.R. Berrone, D.A. Tarzia, L.T. Villa, Asymptotic behavior of a non-classical heat conduction problem for a semi-infinite material, Mathematical Methods in the Applied Sciences 23 (2000) 1161-1177.

\bibitem{BoNaSeTaLibro} J. Bollati,  M.F. Natale, J.A. Semitiel,  D.A. Tarzia,
Approximate solutions to the one-phase Stefan problem with non-linear temperature-dependent thermal conductivity in: J. Hristov – R. Bennacer (Eds.), Heat Conduction: Methods, Applications and Research,  Nova Science Publishers, Inc., 2019, pp.1-20.

\bibitem{BoNaSeTa-ThSci} J. Bollati,  M.F. Natale, J.A. Semitiel, D.A Tarzia, Integral balance methods applied to non-classical Stefan problems, Thermal Science 24 No. 2B (2020) 1229-1241.

\bibitem{BoNaSeTa2020} J. Bollati, M. F. Natale, J. Semitiel, D. A. Tarzia,  Existence and uniqueness of solution for two one-phase Stefan problems with variable thermal coefficients , Nonlinear Analysis: Real World Applications 51 (103001) (2020) 1-11.

\bibitem{BoKh2021} L. Bougoffa, A. Khanfer, Solutions of a non-classical Stefan problem with nonlinear thermal
coefficients and a Robin boundary condition, AIMS Mathematics 6 (6) 6569–6579.

\bibitem{BoTa2000} J. E. Bouillet, D. A. Tarzia, An integral equation for a Stefan problem with many phases and a singular source, Revista de la Uni\'on Matem\'atica Argentina 41 (4) (2000) 1-8.

\bibitem{BoTa2014}	M. Boukrouche, D.A. Tarzia, Global solution to a non-classical heat problem in the space $R^{+}\times R^{n-1}$, Quarterly of Applied Mathematics 72 (2014) 347-361.

\bibitem{BoTa2017}	M. Boukrouche, D.A. Tarzia, Non-classical heat conduction problem with a non local source, Boundary Value Problems 2017 (51) (2017) 1-14.

\bibitem{BoTa2020}	M. Boukrouche, D.A. Tarzia, A heat conduction problem with sources depending on the average of the heat flux on the boundary, Revista de la Uni\'on Matem\'atica Argentina 61 (1) (2020) 87-101.

\bibitem{BrNa3} A.C. Briozzo, M.F. Natale, One-dimensional nonlinear Stefan problems in Storm’s materials, Mathematics, Special Issue on Partial Differential Equations 2 (2014) 1-11.

\bibitem{BrNa2014} A. Briozzo, M. F. Natale, Two Stefan problems for a non-classical heat equation with nonlinear thermal coefficients, Differential and Integral Equations 27 (2014) 1187-1202.

\bibitem{BrNa} A.C. Briozzo,  M.F. Natale, Nonlinear Stefan problem with convective boundary condition in Storm's materials, Z. Angew. Math. Phys.  67(2) (2016) 1-11.

\bibitem{BrNa2019} A. Briozzo, M. F. Natale, Non-classical Stefan problem with nonlinear thermal coefficients and a Robin boundary condition, Nonlinear Analysis: Real World Applications 49 (2019) 159-168.

\bibitem{BrNaTa2007} A. Briozzo, M. F. Natale,  D. A. Tarzia, Explicit solutions for a two-phase unidimensional Lam\'e-Clapeyron-Stefan problem with source terms in both phases, Journal of Mathematical Analysis and Applications 329 (2007) 145-162.

\bibitem{BrTa2006a}	A.C. Briozzo, D.A. Tarzia, Existence and uniqueness of a one-phase Stefan problem for a non-classical heat equation with temperature boundary condition at the fixed face, Electronic Journal of Differential Equations 2006 (21) (2006) 1-16.

\bibitem{BrTa2006b}	A.C. Briozzo, D.A. Tarzia, A one-phase Stefan problem for a non-classical heat equation with a heat flux condition on the fixed face, Applied Mathematics and Computation 182 (2006) 809-819.

\bibitem{BrTa2010a}	A. C. Briozzo, D. A. Tarzia, A Stefan problem for a non-classical heat equation with a convective condition, Applied Mathematics and Computation 217 (2010), 4051-4060.

\bibitem{BrTa2010b}	A. C. Briozzo, D. A. Tarzia, Exact solutions for nonclassical Stefan problems, International Journal of Differential Equations 2010 (868059) (2010) 1-19.

\bibitem{Ca} J.R. Cannon, The One-Dimensional Heat Equation, Addison-Wesley, Menlo Park, CA, 1984.

\bibitem{CaYi1989} J.R. Cannon, H.M. Yin, A class of non-linear non-classical parabolic equations,
J. Diff. Eq. 79 (1989) 266-288.

\bibitem{CaJa} H.S. Carslaw, J.C. Jaeger, Conduction of Heat in Solids, Oxford University Press, London, 1959.

\bibitem{CST}   A.N. Ceretani,  N.N. Salva, D.A.  Tarzia, An exact solution to a Stefan problem with variable thermal conductivity and a Robin boundary condition, Nonlinear Analysis: Real World Applications 40 (2018) 243-259.

\bibitem{CeSaTa2020}	A.N. Ceretani, N.N. Salva, D.A. Tarzia, Auxiliary functions in the study of Stefan-like problems with variable thermal properties, Applied Mathematics Letters 104 (106204) (2020) 1-6.

\bibitem{CeTaVi2015}	A.N. Ceretani, D.A. Tarzia, L.T. Villa, Explicit solutions for a non-classical heat conduction problem for a semi-infinite strip with a non-uniform heat source, Boundary Value Problems 2015 (156) (2015) 1-26.

\bibitem{ChCh} R.M. Cherniha, N.D. Cherniha, Exact solutions of a class of nonlinear boundary value problems with moving boundaries, J. Phys. A: Math. Gen. 26 (1993) L935-L940.

 \bibitem{ChSu74}  S.H. Cho, J.E. Sunderland, Phase-change problems with temperature-dependent thermal conductivity, J. Heat Transfer  96 (1974) 214-217.

\bibitem{Cr} J. Crank, Free and Moving Boundary Problems, Oxford, Clarendon, 1984.


\bibitem{FrJi1988} A. Friedman, L.S. Jiang, Periodic solutions for a thermostat control problem, Comm. Partial Differential Equations 13 (1988) 515-550.

\bibitem{FuMiKe1996} H. Furuya, K. Miyashiba, N. Kenmochi, Asymptotic behavior of solutions of a class of nonlinear evolution equations, J. Differential Equations  62 (1986) 73-94.

\bibitem{GlSp1981} K. Glashoff, J. Sprekels, An application of Glicksberg’s theorem to set-valued integral equations arising in the theory of termostats, SIAM J. Math. Anal. 12 (1981) 477-486.

\bibitem{GlSp1982} K. Glashoff, J. Sprekels, The regulation of temperature by thermostats and set-valued integral equations, J. Integral Equation 4 (1982) 95-112.

\bibitem{Gu} S.C.  Gupta, The Classical Stefan Problem. Basic Concepts, Modelling and Analysis with Quasi-Analytical Solutions and Methods, New Edition, Elsevier, Amsterdam, 2018.

\bibitem{Ke1990} N. Kenmochi, Heat conduction with a class of automatic heat source controls, Pitman Research Notes in Mathematics Series 186 (1990) 471-474.

\bibitem{KePr1988} N. Kenmochi, M. Primicerio, One-dimensional heat conduction with a class of automatic heat source
     controls, IMA J. Appl. Math. 40 (1988) 205-216.

\bibitem{KSR} A. Kumar, A.K. Singh, Rajeev, A moving boundary problem with variable specific heat and thermal conductivity, Journal of King Saud University - Science 32 (2020) 384-389.

\bibitem{Lu} V.J. Lunardini, Heat Transfer With Freezing and Thawing, Elsevier, London, 1991.

\bibitem{Ma} O.D. Makinde, N. Sandeep, T.M. Ajayi,  I. L. Animasaun, Numerical Exploration of Heat Transfer and Lorentz Force Effects on the Flow of MHD Casson Fluid over an Upper Horizontal Surface of a Thermally Stratified Melting Surface of a Paraboloid of Revolution, Int. J. Nonlinear Sci. Simul. 19(2-3) (2018) 93-106.

\bibitem{MeTa1993} J. L. Menaldi, D. A. Tarzia, Generalized Lam\'e-Clapeyron solution for a one-phase source Stefan problem, Comput. Appl. Math. 12 (2) (1993) 123-142.

\bibitem{NaTa} M.F. Natale, D.A. Tarzia, Explicit soltuions to the two-phase Stefan problem for Storm-type materials, J. Phys. A: Math. Gen.  33 (2000) 395-404.

\bibitem{NaTa2003} M.F. Natale, D.A. Tarzia, Explicit solutions to the one-phase Stefan problem with temperature-dependent thermal conductivity and a convective term, Int. J. Engineering Science 41 (2003) 1685-1698.

\bibitem{OlSu87}  D.L.R. Oliver , J.E. Sunderland , A phase-change problem with temperature-dependent thermal conductivity and specific heat, Int. J. Heat Mass Transfer 30 (1987) 2657-2661.

\bibitem{Ro85} C. Rogers, Application of a reciprocal transformation to a two-phase Stefan problem, J. Phys. A: Math. Gen. 18 (1985) L105-L109.

\bibitem{Ro15}  C. Rogers, On a class of reciprocal Stefan moving boundary problems, Z. Angew. Math. Phys. 66 (2015) 2069-2079.

\bibitem{Ro18}  C. Rogers, On Stefan-type moving boundary problems with heterogeneity: canonical reduction via conjugation of reciprocal transformation,  Acta Mech. 230 (2019) 839-850.

\bibitem{SaTaVi2011}	N.N. Salva, D.A. Tarzia, L.T. Villa, An initial-boundary value problem for the one-dimensional non-classical heat equation in a slab, Boundary Value Problems 2011 (4) (2011) 1-17.

\bibitem{Sc1994} E. P. Scott, An analytical solution and sensitivity study of sublimation-dehydration within a porous medium with volumetric heating, J. Heat Transfer 116 (1994) 686-693.

\bibitem{Ta2011} D.A. Tarzia, Explicit and approximated solutions for heat and mass transfer problems with a moving interface, in: M. El-Amin (Ed.), Advanced Topics in Mass Transfer, InTech Open Access Publisher, Rijeka, 2011, pp. 439–484, (Chapter 20).

\bibitem{TaVi1998} D. A.  Tarzia, L. T. Villa, Some nonlinear heat conduction problems for a semi-infinite strip with a non-uniform hear source, Revista de la Uni\'on Matem\'atica Argentina 41 (1998) 99-114.

\bibitem{Vi1986} L. T. Villa,  Problemas de control para una ecuaci\'on unidimensional del calor,  Revista de la Uni\'on Matem\'atica Argentina 32 (1986) 163-169.






\end{thebibliography}

\documentclass[preprint,12pt]{elsarticle}

\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{amsthm}

\usepackage{natbib}

\usepackage{amsmath,amsthm,amsfonts,amssymb,color}
\usepackage[matrix,arrow]{xy}
\usepackage{graphicx}
\usepackage[UKenglish]{babel}
\usepackage{multicol}
\usepackage{appendix}
\usepackage{epstopdf}

\usepackage{caption}
\usepackage{subcaption}

\usepackage{multirow} % para las tablas
\usepackage{float}
\usepackage{graphicx}
\usepackage{longtable}

\newcommand\pder[2][]{\ensuremath{\frac{\partial#1}{\partial#2}}}

\newcommand{\Ste}{\text{Ste}}

\DeclareMathOperator\erf{erf}

 \newtheorem{thm}{Theorem}[section]
 
 \newtheorem{lem}[thm]{Lemma}
 
 \theoremstyle{definition}
 
 \theoremstyle{remark}
 \newtheorem{rem}[thm]{Remark}
 
 \numberwithin{equation}{section}

\begin{document}

\begin{frontmatter}

%% Title, authors and addresses

%% use the tnoteref command within \title for footnotes;
%% use the tnotetext command for theassociated footnote;
%% use the fnref command within \author or \address for footnotes;
%% use the fntext command for theassociated footnote;
%% use the corref command within \author for corresponding author footnotes;
%% use the cortext command for theassociated footnote;
%% use the ead command for the email address,
%% and the form \ead[url] for the home page:
%% \title{Title\tnoteref{label1}}
%% \tnotetext[label1]{}
%% \author{Name\corref{cor1}\fnref{label2}}
%% \ead{email address}
%% \ead[url]{home page}
%%

\title{Explicit solution for non-classical one-phase Stefan problem with variable thermal coefficients and two different heat source terms}

%% use optional labels to link authors explicitly to addresses:
 \author[label1,label2]{Julieta Bollati \corref{cor1}}
 \author[label1]{Mar\'ia F. Natale}
 \author[label1]{Jos\'e A. Semitiel }
 \author[label1,label2]{Domingo A. Tarzia}
\address[label1]{Depto de Matem\'atica, FCE-Universidad Austral,  Paraguay 1950, 2000~Rosario, Argentina}
 \address[label2]{CONICET, Argentina}
\cortext[cor1]{jbollati@austral.edu.ar, (+54) 341 5223000}

\begin{abstract}
A one-phase Stefan problem for a semi-infinite material is investigated for special functional forms of the thermal conductivity and specific heat depending on the temperature of the phase-change material. Using the similarity transformation technique, an explicit solution for these situations are showed. The mathematical analysis is made for two different kinds of heat source terms, and the existence and uniqueness of the solutions are proved.
\end{abstract}

\begin{keyword}
Stefan problem\sep Temperature-dependent thermal coefficients\sep Phase-change material\sep Non-classical heat equation\sep Heat source terms\sep Explicit solution\sep Similarity method.
%% keywords here, in the form: keyword \sep keyword

%% PACS codes here, in the form: \PACS code \sep code

%% MSC codes here, in the form: \MSC code \sep code
%% or \MSC[2008] code \sep code (2000 is the default)

\end{keyword}
\end{frontmatter}

\section{Introduction.}

The phase change problems that contain one or more moving boundaries have attracted growing attention in the last decades due to their wide range of engineering,  industrial applications and natural sciences.  Stefan problems can be modelled as basic phase-change processes where the location of the interface is a priori an unknown function \cite{AlSo,Ca,CaJa,Cr,Gu,Lu,Ta2011}.

The present study provides the existence and uniqueness of solution of the similarity type to a one-phase Stefan fusion problem for a semi-infinite material where it is assumed a Dirichlet condition at the fixed face $x=0$ and it is governed by a non-classical and nonlinear heat equation with temperature-dependent thermal conductivity and specific heat coefficients, and two different kinds of heat source terms.

Non-classical heat conduction problem are considered when the source term is linear or nonlinear depending of the heat flux or the temperature on the boundary of the domain according to the corresponding boundary condition imposed. The non-classical problem are motivated by the modelling of a system of temperature regulation in isotropic media and the source term describes a cooling or a heating effect depending of different types of sources which are related to the evolution of the unknown boundary condition on the boundary of the domain. Problems of this type are related to the thermostat problem \cite{CaYi1989, FrJi1988, FuMiKe1996, GlSp1981, GlSp1982, Ke1990, KePr1988}. For example, we will use mathematical ideas developed for the one-dimensional case in \cite{BeTaVi2000, CeTaVi2015, SaTaVi2011, TaVi1998, Vi1986} and for the n-dimensional case in \cite{BoTa2014, BoTa2017, BoTa2020}. The first paper connecting the non-classical heat equation with a phase-change process (i.e. the Stefan problem) was \cite{BrTa2006a} and after this some other works on the subject were published, for example \cite{BoKh2021, BrNa2014, BrNa2019, BrTa2006b, BrTa2010a}. Moreover, in \cite{BrTa2010b} explicit solutions for the non-classical one-phase Stefan problem was given for cases corresponding to different boundary conditions on the fixed face $x=0$: temperature, heat flux and convective boundary condition.

The mathematical model of the governing phase-change process is described as follows:
\begin{align}
& \rho c(\theta) \pder[\theta]{t}=\frac{\partial}{\partial x} \left(k(\theta)\pder[\theta]{x} \right)-H, & 0<x<s(t), \quad t>0, \label{EcCalor}\\
&  \theta(0,t)=\theta_{_0}>\theta_f &t>0, \label{CondBorde}\\
&  \theta(s(t),t)=\theta_{f}, &t>0, \label{TempCambioFase}\\
&  k_0\pder[\theta]{x}(s(t),t)=-\rho l \dot s(t), &t>0, \label{CondStefan}\\
& s(0)=0,\label{FrontInicial}
\end{align}
where the unknown functions are the temperature $\theta=\theta(x,t)$ and the free boundary $x=s(t)$ separating both phases (the liquid phase at temperature $\theta(x,t)$ and the solid phase at constant temperature $\theta_{f}$ ). The parameters $\rho>0$ (density), $l>0$ (latent heat per unit mass), $\theta_{0}>0$ (temperature imposed at the fixed face $x=0$) and $\theta_f$ (phase change temperature at the free boundary $x=s(t)$) are all known constants.

If the thermal coefficients of the material are temperature-dependent, we have a doubly non-linear free boundary problem.  The functions $k$ and $c$ are defined as:
\begin{align}
&k(\theta)=k_{0}\left(1+\delta\left(\tfrac{\theta-\theta_{f}}{\theta_{0}-\theta_{f}}\right)^{p}\right)\label{k}\\
&c(\theta)=c_{0}\left(1+\delta\left(\tfrac{\theta-\theta_{f}}{\theta_{0}-\theta_{f}}\right)^{p}\right)\label{c},
\end{align}
where $\delta$ and $p$ are given non-negative constants, $k_{0}=k\left(\theta_f\right)$ and $c_{0}=c\left(\theta_f\right)$  are the reference thermal conductivity and the specific heat coefficients, respectively.

Some other models involving temperature-dependent thermal conductivity can also be found in \cite{BoNaSeTaLibro, BoNaSeTa-ThSci,BrNa3, BrNa, CST, CeSaTa2020,ChCh, ChSu74,  KSR, Ma, NaTa, NaTa2003, OlSu87, Ro85, Ro15, Ro18}.

Existence and uniqueness to the problem (\ref{EcCalor})-(\ref{FrontInicial}) with null source term, $H=0$, was developed in \cite{BoNaSeTa2020}.

The control function $H$ represents a heat source term for the nonlinear heat equation. Several applied papers give us the significance of the source term in the interior of the material which can undergo a change of phase \cite{BoTa2000, BrNaTa2007, Sc1994, TaVi1998, Vi1986}.  In this paper we considered two different control functions $H$. The first one is defined as in \cite{BrNaTa2007, BrTa2006b, MeTa1993} and the second one depends on the evolution of the heat flux at the fixed face $x=0$ like in \cite{BrNa2014, BrNa2019}. In this last case, we have a non-classical heat equation as in \cite{TaVi1998, Vi1986}.

We are interested in obtaining a similarity solution to problem (\ref{EcCalor})-(\ref{FrontInicial}) in which the temperature $\theta=\theta(x,t)$ can be written as a function of a single variable.  Through the following change of variables:
\begin{equation}
y(\eta)=\tfrac{\theta(x,t)-\theta_{f}}{\theta_{0}-\theta_{f}}\geq 0  \label{Y}
\end{equation}
with
\begin{equation}
\eta=\tfrac{x}{2a\sqrt{t}},\quad 0<x<s(t),\quad t>0, \label{eta}
\end{equation}
the phase front moves as
\begin{equation}
s(t)=2a\lambda\sqrt{t} \label{freeboundary}
\end{equation}
where $a^{2}=\frac{k_{0}}{\rho c_{0}}$ (thermal diffusivity) and $\lambda>0$ is a positive parameter to be determined.

The plan of this paper is the following. In Section 2, we prove the existence and uniqueness of solution to the problem (\ref{EcCalor})-(\ref{FrontInicial}) considering the control function given by \cite{Sc1994}:
\begin{equation}
H=H_1(x,t)=\frac{\rho l}{t}\beta\left(\frac{x}{2a\sqrt{t}}\right) \label{H1},
\end{equation}
where $\beta=\beta(\eta)$ in a function with appropriate regularity properties \cite{BoTa2000, BrNaTa2007, MeTa1993, Sc1994}.  Moreover, a particular case where $\beta$ is of exponential type given by
\begin{equation}
\beta(\eta)=\dfrac{1}{2}\exp(-\eta^2),
\end{equation}
is also studied in details. This type of heat source term is important through the use of microwave energy following \cite{Sc1994}.

Finally,  in Section 3, we prove existence and uniqueness of solution to the problem (\ref{EcCalor})-(\ref{FrontInicial}) considering the control function given by
\begin{equation}
H=H_2(t)=\frac{\lambda_0 }{\sqrt{t}}\frac{\partial T}{\partial x}(0,t)
\end{equation}
that can be thought of  by modelling of a system of temperature regulation in isotropic mediums \cite{BrNa2014, BrNa2019, BrTa2006b} with nonuniform source term, which provides a cooling or helting effect depending upon the properties of $H_2$ related to the heat flux (or the temperature in other cases) at the fixed face boundary $x=0$.

\section{Free boundary problem when the heat source term is of a similarity type}
We consider now the control function $H$ given by (\ref{H1}).

% $$H=H_1(x,t)=\frac{\rho l}{t}\beta\left(\frac{x}{2a\sqrt{t}}\right)$$

\subsection{General case}
We consider the hypothesis:
\\
$H_\beta$: $\beta=\beta(\eta)$ is an integrable function in $\left(0,\epsilon\right)$ for all $\epsilon >0$ and $\beta(\eta)\exp\left(\eta^2\right)$ is an integrable function in $\left(M,+\infty\right)$ for all $M>0$ \label{Hbeta}.

It is easy to see that the Stefan problem $\mathrm{(}$\ref{EcCalor}$\mathrm{)}$-(\ref{FrontInicial}) has a similarity solution $(\theta,s)$ given by:
\begin{align}
&\theta(x,t)=\left(\theta_{0}-\theta_{f}\right)y\left(\tfrac{x}{2a\sqrt{t}}\right)+\theta_{f},\quad  0<x<s(t), \quad t>0,\label{T1} \\
&s(t)=2a\lambda\sqrt{t},\quad\quad t>0\label{s1}
\end{align}
if and only if the function $y$ (defined by (\ref{Y})) and the parameter $\lambda>0$ (defined in (\ref{freeboundary})) satisfy the following ordinary differential problem:
\begin{align}
&2\eta(1+\delta y^p(\eta))y'(\eta)+[(1+\delta y^p(\eta))y'(\eta)]'=\tfrac{4}{\Ste}\beta(\eta), \quad &0<\eta<\lambda, \label{y1}\\
&y(0)=1,\label{cond01}\\
&y(\lambda)=0, \label{condlambda1}\\
&y'(\lambda)=-\tfrac{2\lambda}{\Ste} \label{eclambda1}
\end{align}
where $\delta > 0$, $p> 0$ and $\Ste=\tfrac{c_{0}(T_{0}-T_f)}{l}>0$ is the Stefan number.

\begin{lem}\label{ProbAux1}
Fixed $p > 0$, $\delta > 0$, $\lambda>0$, $y\in C^{\infty}[0,\lambda]$, $y\geq 0$, and $\beta=\beta(\eta)$  verifies the hypothesis $H_\beta$.

%an integrable function in $\left(0,\epsilon\right)$ for all $\epsilon >0$ and $\beta(\eta)\exp\left(\eta^2\right)$ %an integrable function in $\left(M,+\infty\right)$ for all $M>0$.

Then, $(y,\lambda)$ is a solution to the ordinary differential equation (\ref{y1})-(\ref{eclambda1}) if and only if  $\lambda>0$ is a solution to the equation:
\begin{align}\label{7-1}
&\tfrac{\sqrt{\pi}}{{\Ste}} x \erf(x) \exp(x^2) +\tfrac{2\sqrt{\pi}}{\Ste} \int_{0}^{x} \exp(\xi^2)\erf(\xi)\beta(\xi)\, d\xi =1+\tfrac{\delta}{p+1}, \quad x>0,
\end{align}
and the function $y=y(\eta)$ satisfies the functional equation:
\begin{align}\label{8-1}
y(\eta)\left(1+\tfrac{\delta}{p+1}y^p(\eta)\right)&=1+\tfrac{\delta}{p+1}-\tfrac{\sqrt{\pi} \erf(\eta)}{\mathrm{Ste}}\left(2\int_{0}^{\lambda} \beta(\xi) \exp(\xi^2)\, d\xi - \lambda \exp(\lambda^2)\right)+\nonumber\\
&+\tfrac{2\sqrt{\pi}}{\Ste}\left( \int_{0}^{\eta} \beta(\xi) \exp(\xi^2) \left( \erf(\eta) - \erf(\xi)\right)\, d\xi\right), \ 0 \leq \eta \leq \lambda
\end{align}
 \end{lem}

\begin{proof}
Let $(y,\lambda)$ be a solution to (\ref{y1})-(\ref{eclambda1}). We define the function:
\begin{equation}
v(\eta)=\left(1+\delta y^{p}(\eta) \right) y'(\eta). \label{v-eta-1}
\end{equation}

Taking into account  (\ref{y1}) and the condition  (\ref{cond01})  the function $v$ can be rewrite as
\begin{equation}
v(\eta)=\exp{(-\eta^2)}\left(\tfrac{4}{\Ste}\int_{0}^{\eta} \beta(\xi)\exp(\xi^2)\, d\xi+ (1+\delta)y'(0)\right). \label{v-eta1-2}
\end{equation}

From (\ref{v-eta-1}) and (\ref{v-eta1-2}), we obtain
\begin{equation} \label{ec2-11-1}
\left(1+\delta y^{p}(\eta) \right) y'(\eta)=\exp{(-\eta^2)}\left(\tfrac{4}{\Ste}\int_{0}^{\eta} \beta(\xi)\exp(\xi^2)\, d\xi+ (1+\delta)y'(0)\right).
\end{equation}

Taking $\eta=\lambda$ in the above equation, using (\ref{condlambda1}) and (\ref{eclambda1}), we obtain:
\begin{equation}
y'(0)=-\tfrac{4}{\Ste(1+\delta)}\left(2\int_{0}^{\lambda} \beta(\xi)\exp(\xi^2)\, d\xi+\lambda \exp{(\lambda^2)}\right) \label{yprima01}
\end{equation}

Integrating equation (\ref{ec2-11-1}) in the domain $(0,\eta)$  and by virtue of (\ref{cond01}), we obtain:

\begin{align}
y(\eta)\left(1+\tfrac{\delta}{p+1}y^p(\eta)\right)&=1+\tfrac{\delta}{p+1}+(1+\delta)y'(0)\tfrac{\sqrt{\pi}}{2}\erf(\eta)+\nonumber\\
&+\tfrac{4}{\Ste}\int_{0}^{\eta} \int_{\xi}^{\eta} \beta(\xi)\exp(-z^2)\exp(\xi^2)\, dz\ d\xi
\end{align}
Given that $\int_{0}^{\eta} \int_{\xi}^{\eta} \beta(\xi)\exp(-z^2)\exp(\xi^2)\, dz\ d\xi =\tfrac{\sqrt{\pi}}{2} \int_{0}^{\eta} \left( \erf(\eta)-\erf(\xi)\right)\beta(\xi)\exp(\xi^2)\, d\xi$ and from (\ref{yprima01}), we obtain that $y=y(\eta)$ is a solution to  (\ref{8-1}).

\medskip

Taking $\eta=\lambda$ in equation (\ref{8-1}) and using (\ref{condlambda1}), we conclude that $\lambda>0$ is a solution to equation (\ref{7-1}).

\medskip

Reciprocally, if $(y,\lambda)$ is a solution to  (\ref{7-1})-(\ref{8-1}), then
\begin{align*}
y(\eta)&=1+\tfrac{\delta}{p+1}-\tfrac{\delta}{p+1}y^{p+1}(\eta)-\tfrac{\sqrt{\pi} \erf(\eta)}{\mathrm{Ste}}\left(2\int_{0}^{\lambda} \beta(\xi) \exp(\xi^2)\, d\xi - \lambda \exp(\lambda^2)\right)+\nonumber\\
&+\tfrac{2\sqrt{\pi}}{\Ste}\left( \int_{0}^{\eta} \beta(\xi) \exp(\xi^2) \left( \erf(\eta) - \erf(\xi)\right)\, d\xi\right),
\end{align*}

and it follows immediately that  $(y,\lambda)$ is a solution to (\ref{y1})-(\ref{eclambda1}).

\end{proof}

\begin{lem}\label{ProbAux11}
If $p > 0$, $\delta > 0$, $\beta=\beta(\eta)\geq 0$  verifies the hypothesis $H_\beta$, then there exists a unique solution $(y,\lambda)$ to the functional problem defined by (\ref{7-1})-(\ref{8-1}) with $\lambda>0$, $y\in C^{\infty}[0,\lambda]$ and $y\geq 0$.
\end{lem}
%an integrable function in $\left(0,\epsilon\right)$ for all $\epsilon >0$ and $\beta(\eta)\exp\left(\eta^2\right)$ %an integrable function in $\left(M,+\infty\right)$ for all $M>0$,

\begin{proof}
For $x>0$, we define the function given by $$\varphi_1(x)=\tfrac{\sqrt{\pi}}{{\Ste}} x \erf(x) \exp(x^2) +\tfrac{2\sqrt{\pi}}{\Ste} \int_{0}^{x} \exp(\xi^2)\erf(\xi)\beta(\xi)\, d\xi .$$

It is easy to see that $\varphi_1(0)=0$, $\varphi_1(+\infty)=+\infty$ and $\varphi_1$ is an increasing function. Then, there exists a unique  $\lambda>0$ solution to equation  (\ref{7-1}): $$\varphi_1(x)= 1+\tfrac{\delta}{p+1}.$$

Let $\Phi$ be the function defined by 
\begin{equation}\label{funcionphi}
 \Phi(x)=x+\tfrac{\delta}{p+1}x^{p+1}, \quad 0\leq x \leq1.
\end{equation}
  Clearly,  $\Phi(0)=0$, $\Phi(1)=1+\tfrac{\delta}{p+1}$ and $\Phi$ in an increasing function then, there exists the inverse function $\Phi^{-1}:[0,1+\tfrac{\delta}{p+1}] \to [0,1]$.
\medskip

For the unique value $\lambda>0$ obtained in (\ref{7-1}), we can define the function
\begin{align*}
\Psi_1(x)&=1+\tfrac{\delta}{p+1}-\tfrac{\sqrt{\pi} \erf(x)}{\mathrm{Ste}}\left(2\int_{0}^{\lambda} \beta(\xi) \exp(\xi^2)\, d\xi - \lambda \exp(\lambda^2)\right)+\nonumber\\
&+\tfrac{2\sqrt{\pi}}{\Ste}\left( \int_{0}^{x} \beta(\xi) \exp(\xi^2) \left( \erf(x) - \erf(\xi)\right)\, d\xi\right), \quad x \in \left[0,\lambda\right].
\end{align*}
Then $\Psi_1(0)=1+\tfrac{\delta}{p+1}$,  $\Psi_1(\lambda)=0$ and $\Psi_1$ is a decreasing function. Furthermore, $\Psi_1(x) \in \left[0,1+\tfrac{\delta}{p+1}\right]$ for all $x \in \left[0,\lambda\right]$.

\medskip

We conclude that there exists a unique function $y\in C^{\infty}[0,\lambda]$ solution to the equation  $$\Phi\left(y(\eta)\right)=\Psi_1(\eta),$$ given by
\begin{equation}
y(\eta)=\Phi^{-1}\left(\Psi_1(\eta)\right), \quad 0\leq \eta \leq \lambda.
\end{equation}
\end{proof}

From the above lemmas we are able to claim the following result:

\begin{thm} \label{ExistenciaDirichlet-1}
The Stefan problem governed by $($\ref{EcCalor}$)$-$($\ref{FrontInicial}$)$ has a unique similarity type solution given by $($\ref{T1}$)$-$($\ref{s1}$)$ where $(y,\lambda)$ is the unique solution to the functional problem $($\ref{7-1}$)$-$($\ref{8-1}$)$.
\end{thm}

\begin{rem} \label{2.4-1}
On one hand we have that $\Phi$ is an increasing function with $\Phi(0)=0$ and $\Phi(1)=1+\frac{\delta}{p+1}$.  On the other hand, $\Psi_1$ is a decreasing function with  $\Psi_1(0)=1+\frac{\delta}{p+1}$ and $\Psi_1(\lambda)=0$. Then it follows that $0\leq y(\eta)\leq 1 $, for $0\leq\eta\leq\lambda$.

In virtue of this and Theorem \ref{ExistenciaDirichlet-1} we have that
\begin{equation*}
\theta_f<\theta(x,t)<\theta_0,\qquad \qquad 0<x<s(t),\quad t>0.
\end{equation*}
\end{rem}

\subsection{Particular case}
If we consider the particular case where $\beta$ is of exponential type given by
$$\beta(\eta)=\dfrac{1}{2}\exp(-\eta^2)$$

\begin{thm} \label{ExistenciaDirichlet-11}
If $p > 0$, $\delta > 0$, the Stefan problem governed by $($\ref{EcCalor}$)$-$($\ref{FrontInicial}$)$ has a unique similarity type solution $(y,\lambda)$
with $\lambda>0$, $y\in C^{\infty}[0,\lambda]$ and $y\geq 0$, where  $\lambda>0$ is the unique solution to the equation:
\begin{align}
\tfrac{1}{\Ste}\left(1-\exp(-x^2)\right)+\tfrac{\sqrt{\pi}}{\Ste}x  ~\erf(x)\left(\exp(x^2)-1\right)=1+\tfrac{\delta}{p+1},\qquad \qquad x>0, \label{7-1-1}
\end{align}
and the function $y=y(\eta)$ satisfies the equation
\begin{align} \label{8-1-1}
y(\eta)\left(1+\tfrac{\delta}{p+1}y^p(\eta)\right)=1+\tfrac{\delta}{p+1}+\tfrac{1}{\Ste}\left( \exp(-\eta^2)-1\right)+   \tfrac{\sqrt{\pi}}{\mathrm{Ste}} \;\lambda ~ \erf(\eta)\left(1-\exp(\lambda^2)\right),\nonumber\\
0 \leq \eta \leq \lambda.
\end{align}

\end{thm}

\begin{rem}
For the particular case $p=1$, the unique function $y=y(\eta)$ solution to equation (\ref{8-1-1}) is given by
\begin{equation*}
y(\eta)=\tfrac{1}{\delta} \left[\sqrt{1+4\left(1+\tfrac{\delta}{2}\right)+\tfrac{1}{\Ste}\left(\exp(-\eta^2-1\right)+\tfrac{\sqrt{\pi}}{\Ste} \lambda \erf(\eta) \left(1-\exp(\lambda^2)\right)}-1 \right],
\end{equation*}
where $\lambda>0$ is the unique solution to equation (\ref{7-1-1}) for $p=1$.
\end{rem}

\section{Free boundary problem with a heat source that depends on the evolution of the heat flux at the fixed face $x=0$}
If we consider the control function $H$ depends on the evolution of the heat flux at the fixed face $x=0$, that is $$H=H_2(W(t),t)=\frac{\lambda_0 }{\sqrt{t}}\frac{\partial T}{\partial x}(0,t),$$
where $\lambda_0>0$, it is easy to see that the Stefan problem $\mathrm{(}$\ref{EcCalor}$\mathrm{)}$-(\ref{FrontInicial}) has a similarity solution $(\theta,s)$ given by:
\begin{align}
&\theta(x,t)=\left(\theta_{0}-\theta_{f}\right)y\left(\tfrac{x}{2a\sqrt{t}}\right)+\theta_{f},\quad  0<x<s(t), \quad t>0,\label{T} \\
&s(t)=2a\lambda\sqrt{t},\quad\quad t>0\label{s}
\end{align}
if and only if the function $y$ and the parameter $\lambda>0$ satisfy the following ordinary differential problem:
\begin{align}
&2\eta(1+\delta y^p(\eta))y'(\eta)+[(1+\delta y^p(\eta))y'(\eta)]'=Ay'(0), \quad &0<\eta<\lambda, \label{y}\\
&y(0)=1,\label{cond0}\\
&y(\lambda)=0, \label{condlambda}\\
&y'(\lambda)=-\tfrac{2\lambda}{\Ste} \label{eclambda}
\end{align}
where $\delta > 0$, $p> 0$, $A=\frac{2\lambda_0}{\rho c_0 a}$ and $\Ste=\tfrac{c_{0}(T_{0}-T_f)}{l}>0$ is the Stefan number.

\begin{lem}\label{ProbAux}
Fixed $p > 0$, $\delta > 0$, $\lambda>0$, $y\in C^{\infty}[0,\lambda]$ and $y\geq 0$. Then, $(y,\lambda)$ is a solution to the ordinary differential equation (\ref{y})-(\ref{eclambda}) if and only if  $\lambda$ is a solution to the equation:
\begin{align}\label{7}
&\tfrac{\sqrt{\pi}x \exp{(x^2)}}{\Ste \left(A \int_{0}^{x}  \exp(z^2)\, dz +1 +\delta \right) } \left(A\int_{0}^{x} \exp(z^2)\left(\erf(x)-\erf(z)\right)\, dz+(1+\delta)\erf(x)\right)=\nonumber\\
& \qquad \qquad \qquad \qquad \qquad \qquad \qquad \qquad \qquad \qquad \qquad \qquad =1+\tfrac{\delta}{p+1}, \quad x>0,
\end{align}
and the function $y=y(\eta)$ satisfies the equation:
\begin{align}\label{8}
y(\eta)\left(1+\tfrac{\delta}{p+1}y^p(\eta)\right)&=1+\tfrac{\delta}{p+1}- \tfrac{\sqrt{\pi} \lambda\exp(\lambda^2)}{\Ste\left(A\int_{0}^{\lambda} \exp(z^2)\, dz+1+\delta\right)}    \left(A\int_{0}^{\eta} \exp(z^2)\left(\erf(\eta)+\right.\right.\nonumber\\
&\left.\left.-\erf(z)\right)\, dz + (1+\delta)\erf(\eta)\right), \qquad \qquad \qquad 0\leq \eta\leq \lambda.
\end{align}
 \end{lem}

\begin{proof}
Let $(y,\lambda)$ be a solution to (\ref{y})-(\ref{eclambda}). We define the function:
\begin{equation}
v(\eta)=\left(1+\delta y^{p}(\eta) \right) y'(\eta). \label{v-eta}
\end{equation}

Taking into account  (\ref{y}) and the condition  (\ref{cond0})  the function $v$ can rewrite as
\begin{equation}
v(\eta)=\exp{(-\eta^2)}y'(0)\left(A\int_{0}^{\eta} \exp(z^2)\, dz+ 1+\delta\right). \label{v-eta1}
\end{equation}

From (\ref{v-eta}) and (\ref{v-eta1}), we obtain
\begin{equation} \label{ec2-11}
\left(1+\delta y^{p}(\eta) \right) y'(\eta)=\exp{(-\eta^2)}y'(0)\left(A\int_{0}^{\eta} \exp(z^2)\, dz+ 1+\delta\right).
\end{equation}

Taking $\eta=\lambda$ in the above equation, using (\ref{condlambda}) and (\ref{eclambda}), we obtain:
\begin{equation}
y'(0)=-\frac{2\lambda\exp(\lambda^2)}{\Ste\left(A\int_{0}^{\lambda} \exp(z^2)\, dz+1+\delta\right)} \label{yprima0}
\end{equation}

Integrating into $(0,\eta)$ equation (\ref{ec2-11}) and by virtue of (\ref{cond0}), we obtain:

\begin{align}
y(\eta)\left(1+\tfrac{\delta}{p+1}y^p(\eta)\right)&=1+\tfrac{\delta}{p+1}+(1+\delta)y'(0)\tfrac{\sqrt{\pi}}{2}\erf(\eta)+\nonumber\\
&+Ay'(0)\int_{0}^{\eta} \int_{z}^{\eta} \exp(-z^2)\exp(\xi^2)\, dz\ d\xi
\end{align}
Given that $\int_{0}^{\eta} \int_{z}^{\eta}\exp(-z^2)\exp(\xi^2)\, dz\ d\xi =\tfrac{\sqrt{\pi}}{2} \int_{0}^{\eta} \left( \erf(\eta)-\erf(z)\right)\exp(z^2)\, dz$ and from (\ref{yprima0}), we obtain that $y=y(\eta)$ is a solution to  (\ref{8}).

\medskip

Taking $\eta=\lambda$ in equation (\ref{8}) and using (\ref{condlambda}), we conclude that $\lambda>0$ is a solution to equation (\ref{7}).

\medskip

Reciprocally, if $(y,\lambda)$ is a solution to  (\ref{7})-(\ref{8}),
\begin{align*}
y(\eta)&=1+\tfrac{\delta}{p+1}-\tfrac{\delta}{p+1}y^{p+1}(\eta)- \tfrac{\sqrt{\pi} \lambda\exp(\lambda^2)}{\Ste\left(A\int_{0}^{\lambda} \exp(z^2)\, dz+1+\delta\right)}    \left(A\int_{0}^{\eta} \exp(z^2)\left(\erf(\eta)+\right.\right.\nonumber\\
&\left.\left.-\erf(z)\right)\, dz + (1+\delta)\erf(\eta)\right),
\end{align*}

and it follows immediately that  $(y,\lambda)$ is a solution to (\ref{y})-(\ref{eclambda}).

\end{proof}

\begin{lem}\label{ProbAux}
If $p > 0$ y $\delta > 0$, then there exists a unique solution $(y,\lambda)$ to the functional problem defined by (\ref{7})-(\ref{8}) with $\lambda>0$, $y\in C^{\infty}[0,\lambda]$ and $y\geq 0$.
\end{lem}
\begin{proof}
For $x>0$, we define the function given by $$\varphi_2(x)=\tfrac{\sqrt{\pi}x \exp{(x^2)}}{\Ste \left(A \int_{0}^{x}  \exp(z^2)\, dz +1 +\delta \right) } \left(A\int_{0}^{x} \exp(z^2)\left(\erf(x)-\erf(z)\right)\, dz+(1+\delta)\erf(x)\right).$$

It is easy to see that $\varphi_2(0)=0$, $\varphi_2(+\infty)=+\infty$ and $\varphi_2$ is an increasing function. Then, there exists a unique  $\lambda>0$ solution to equation  (\ref{7}): $$\varphi_2(x)= 1+\tfrac{\delta}{p+1}.$$

Let $\Phi$ be the function given by \eqref{funcionphi}.  Clearly,  $\Phi(0)=0$,  $\Phi(1)=1+\tfrac{\delta}{p+1}$ and $\Phi$ in an increasing function then, there exists the function $\Phi^{-1}:[0,1+\tfrac{\delta}{p+1}] \to [0,1]$.
\medskip

For the unique value $\lambda>0$ obtained in (\ref{7}), let the function
\begin{align*}
\Psi_2(x)&=1+\tfrac{\delta}{p+1}- \tfrac{\sqrt{\pi} \lambda\exp(\lambda^2)}{\Ste\left(A\int_{0}^{\lambda} \exp(z^2)\, dz+1+\delta\right)}\left(A\int_{0}^{x} \exp(z^2)\left(\erf(x)-\erf(z)\right)\, dz+\right.\\
&\left.+(1+\delta)\erf(x)\right), \quad x \in \left[0,\lambda\right].
\end{align*}
Then $\Psi_2(0)=1+\tfrac{\delta}{p+1}$,  $\Psi_2(\lambda)=0$ and $\Psi_2$ is a decreasing function.  Furthermore,  $\Psi_2(x) \in \left[0,1+\tfrac{\delta}{p+1}\right]$ for all $x \in \left[0,\lambda\right]$.

\medskip

We conclude that there exists a unique function $y\in C^{\infty}[0,\lambda]$ solution to the equation  $$\Phi\left(y(\eta)\right)=\Psi_2(\eta),$$ given by
\begin{equation}
y(\eta)=\Phi^{-1}\left(\Psi_2(\eta)\right), \quad 0 \leq \eta \leq \lambda.
\end{equation}
\end{proof}

From the above lemmas we are able to claim the following result:

\begin{thm} \label{ExistenciaDirichlet}
The Stefan problem governed by $($\ref{EcCalor}$)$-$($\ref{FrontInicial}$)$ has a unique similarity type solution given by $($\ref{T}$)$-$($\ref{s}$)$ where $(y,\lambda)$ is the unique solution to the functional problem $($\ref{7}$)$-$($\ref{8}$)$.
\end{thm}

\begin{rem} \label{2.4}
On one hand we have that $\Phi$ is an increasing function with $\Phi(0)=0$ and $\Phi(1)=1+\frac{\delta}{p+1}$.  On the other hand, $\Psi_2$ is a decreasing function with  $\Psi_2(0)=1+\frac{\delta}{p+1}$ and $\Psi_2(\lambda)=0$. Then it follows that $0\leq y(\eta)\leq 1 $, for $0 \leq \eta \leq \lambda$.

From this and Theorem \ref{ExistenciaDirichlet} we have that
\begin{equation*}
\theta_f<\theta(x,t)<\theta_0,\qquad \qquad 0<x<s(t),\quad t>0.
\end{equation*}
\end{rem}

\section{Conclusions}
One dimensional non-classical Stefan problems with temperature dependent thermal coefficients and a Dirichlet type condition at fixed face $x=0$ for a semi-infinite phase-change material were considered.  Existence and uniqueness of solution were obtained by using the similarity method and explicit solutions were found.
% ------------------------------------------------------------------------

\section*{Acknowledgement}
The present work has been partially sponsored by the Project PIP No 0275 from CONICET - Universidad Austral, Rosario, Argentina, and by the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement No. 823731 CONMECH.

\end{document}